What is the circuit (amps) for a parallel circuit containing a 12 V potential and three resistors R1= 3 ohms, R2=2 ohms, and R3= 1 ohm?


1
Expert's answer
2021-05-09T15:43:03-0400

Net resistance-


"\dfrac{1}{R}=\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3}"


"=\dfrac{1}{3}+\dfrac{1}{2}+\dfrac{1}{1}

 \\[9pt] =\dfrac{2+3+6}{6}\\[9pt]

 =\dfrac{11}{6}"


So "R=\dfrac{6}{11} \Omega"


Current in the circuit "=\dfrac{V}{R}=\dfrac{12}{\frac{6}{11}}=22 amps"